OpenStudy (anonymous):

Identify the asymptote: f(x) = -2/ x + 3

OpenStudy (anonymous):

there would be a vertical asymptote at the value(s) of x which make the function not exist. For example, when plugging x into f(x) would make you have to divide by zero (which is not allowed)

sam (.sam.):

\[f(x)=-\frac{2}{x+3} ~~~~or~~~~-\frac{2}{x}+3\]

OpenStudy (jhannybean):

A vertical asymptote occurs when the denominator = 0 or dne, so set \(\large \color{orchid}{x+3}=\color{green}{0}\)